OpenStudy (anonymous):

Two forces with magnitudes of 150 and 75 pounds act on an object at angles of 30° and 150° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

sum of x-components= 150cos(30) + 75cos(150) = 129.90 + (-64.95) = 64.95 sum of y-components= 150sin(30) + 75sin(150) = 75+ 37.5 = 112.5 magnitude = \[\sqrt{x ^{2}+y ^{2}}\] = \[\sqrt{(64.95)^{2}+ (112.5)^{2}}\] = 129.90 direction = arctan y/x = arctan 112.5/64.95 = 60.00 final answer = 129.90 pounds,60 degrees

OpenStudy (anonymous):

Thank you! @conan.195114